CaO can be used as a drying agent. One such application occurs when water is added to dry concrete or cement. The reaction that occurs is

CaO(s)+H2O(l)⇌Ca(OH)2(s)

The product is commonly called slaked lime.

Assuming the commonly used standard-state temperature of 25∘C, calculate ΔSuniv for this reaction using table from the table below.

Substance S
[J/(K⋅mol)]
ΔH∘f
(kJ/mol)
CaO(s) 39.9 −635.1
H2O(l) 69.9 −285.8
Ca(OH)2(s) 83.4 −986.1
Answer #1
Concepts and reason

The reaction is as follows:

CaO(s)+ H2O(Ca(OH), (s)

Calculate the entropy change of the universe by using the following formula.

ΔSUniverse=ΔSSystem+ΔSSurroundings\Delta {S_{{\rm{Universe}}}} = {\rm{ }}\Delta {S_{{\rm{System}}}} + \Delta {S_{{\rm{Surroundings}}}}

Calculate the ΔSSurroundings\Delta {S_{{\rm{Surroundings}}}} as shown below.

ΔSSurroundings=ΔHorxnT\Delta {S_{{\rm{Surroundings}}}} = - \frac{{\Delta {H^{\rm{o}}}_{{\rm{rxn}}}}}{T}

Here, ΔHorxn\Delta {H^{\rm{o}}}_{{\rm{rxn}}} is enthalpy change of the reaction, T is temperature.

Calculate the entropy change for the reaction (system) by using the following formula.

ΔSoreaction=npSof(products)nrSof(reactants)\Delta {S^{\rm{o}}}_{{\rm{reaction}}} = {\rm{ }}\sum {n_p}{S^{\rm{o}}}_{\rm{f}}{\rm{(products)}} - \sum {n_r}{S^{\rm{o}}}_{\rm{f}}{\rm{(reactants)}}

Change of enthalpy during the formation of substance from its constituent elements is called as enthalpy of formation.

Calculate the enthalpy change for the reaction by using the following formula.

ΔHoreaction=npΔHof(products)nrΔHof(reactants)\Delta {H^{\rm{o}}}_{{\rm{reaction}}} = {\rm{ }}\sum {n_p}\Delta {H^{\rm{o}}}_{\rm{f}}{\rm{(products)}} - \sum {n_r}\Delta {H^{\rm{o}}}_{\rm{f}}{\rm{(reactants)}}

Fundamentals

Entropy is a thermodynamic quantity, it is a measure of randomness of a system.

Change in the enthalpy when one mole of a substance is formed from its pure elements under standard conditions is called as standard enthalpy of formation. The standard conditions are 1atm pressure and 298 K temperature.

Write the expansion of enthalpy formation for the chemical equation as shown below.

ΔSorxn=[nCa(OH)2(s)×SofCa(OH)2(s)][nCaO(s)×SofCaO(s)+nH2O(l)×SofH2O(l)]ΔSorxn=[1mol×83.4JKmol][(1mol×39.9JKmol)+(1mol×69.9JKmol)]\begin{array}{c}\\\Delta {S^{\rm{o}}}_{{\rm{rxn}}} = \left[ {{n_{{\rm{Ca}}{{\left( {{\rm{OH}}} \right)}_{\rm{2}}}\left( {\rm{s}} \right)}} \times {S^{\rm{o}}}_{\rm{f}}{\rm{Ca}}{{\left( {{\rm{OH}}} \right)}_{\rm{2}}}\left( {\rm{s}} \right)} \right] - \left[ {{n_{{\rm{CaO}}\left( {\rm{s}} \right)}} \times {S^{\rm{o}}}_{\rm{f}}{\rm{CaO}}\left( {\rm{s}} \right) + {n_{{{\rm{H}}_{\rm{2}}}{\rm{O}}\left( {\rm{l}} \right)}} \times {S^{\rm{o}}}_{\rm{f}}{{\rm{H}}_{\rm{2}}}{\rm{O}}\left( {\rm{l}} \right)} \right]\\\\\Delta {S^{\rm{o}}}_{{\rm{rxn}}} = \left[ {1{\rm{ mol}} \times 83.4{\rm{ }}\frac{{\rm{J}}}{{{\rm{K}} \cdot {\rm{mol}}}}} \right] - \left[ {\left( {1{\rm{ mol}} \times 39.9{\rm{ }}\frac{{\rm{J}}}{{{\rm{K}} \cdot {\rm{mol}}}}} \right) + \left( {{\rm{1 mol}} \times {\rm{69}}{\rm{.9 }}\frac{{\rm{J}}}{{{\rm{K}} \cdot {\rm{mol}}}}} \right)} \right]\\\end{array}

ΔSorxn=[1mol×83.4JKmol][(1mol×39.9JKmol)+(1mol×69.9JKmol)]ΔSorxn=83.4JK[39.9JK+69.9JK]ΔSorxn=26.4JK\begin{array}{c}\\\Delta {S^{\rm{o}}}_{{\rm{rxn}}} = \left[ {1{\rm{ }}\cancel{{{\rm{mol}}}} \times 83.4{\rm{ }}\frac{{\rm{J}}}{{{\rm{K}} \cdot \cancel{{{\rm{mol}}}}}}} \right] - \left[ {\left( {1{\rm{ }}\cancel{{{\rm{mol}}}} \times 39.9{\rm{ }}\frac{{\rm{J}}}{{{\rm{K}} \cdot \cancel{{{\rm{mol}}}}}}} \right) + \left( {{\rm{1 }}\cancel{{{\rm{mol}}}} \times {\rm{69}}{\rm{.9 }}\frac{{\rm{J}}}{{{\rm{K}} \cdot \cancel{{{\rm{mol}}}}}}} \right)} \right]\\\\\Delta {S^{\rm{o}}}_{{\rm{rxn}}} = 83.4{\rm{ }}\frac{{\rm{J}}}{{\rm{K}}} - \left[ {39.9{\rm{ }}\frac{{\rm{J}}}{{\rm{K}}} + 69.9{\rm{ }}\frac{{\rm{J}}}{{\rm{K}}}} \right]\\\\\Delta {S^{\rm{o}}}_{{\rm{rxn}}} = - {\rm{26}}{\rm{.4 }}\frac{{\rm{J}}}{{\rm{K}}}\\\end{array}

Thus, ΔSSystem\Delta {S_{{\rm{System}}}} of the reaction is 26.4JK - {\rm{26}}{\rm{.4 }}\frac{{\rm{J}}}{{\rm{K}}} .

Calculate the enthalpy change for reaction as shown below.

ΔHorxn=[nCa(OH)2(s)×ΔHofCa(OH)2(s)][nCaO(s)×ΔHofCaO(s)+nH2O(l)×ΔHofH2O(l)]ΔHorxn=1mol×(986.1kJmol)[1mol×(635.1kJmol)+1mol×(285.8kJmol)]\begin{array}{c}\\\Delta {H^{\rm{o}}}_{{\rm{rxn}}} = {\rm{ }}\left[ {{n_{{\rm{Ca}}{{\left( {{\rm{OH}}} \right)}_{\rm{2}}}\left( {\rm{s}} \right)}} \times \Delta {H^{\rm{o}}}_{\rm{f}}{\rm{Ca}}{{\left( {{\rm{OH}}} \right)}_{\rm{2}}}\left( {\rm{s}} \right)} \right] - \left[ {{n_{{\rm{CaO}}\left( {\rm{s}} \right)}} \times \Delta {H^{\rm{o}}}_{\rm{f}}{\rm{CaO}}\left( {\rm{s}} \right) + {n_{{{\rm{H}}_{\rm{2}}}{\rm{O}}\left( {\rm{l}} \right)}} \times \Delta {H^{\rm{o}}}_{\rm{f}}{{\rm{H}}_{\rm{2}}}{\rm{O}}\left( {\rm{l}} \right)} \right]\\\\\Delta {H^{\rm{o}}}_{{\rm{rxn}}} = 1{\rm{ mol}} \times \left( { - 986.1{\rm{ }}\frac{{{\rm{kJ}}}}{{{\rm{mol}}}}} \right) - \left[ {1{\rm{ mol}} \times \left( { - 635.1{\rm{ }}\frac{{{\rm{kJ}}}}{{{\rm{mol}}}}} \right) + 1{\rm{ mol}} \times \left( { - {\rm{285}}{\rm{.8 }}\frac{{{\rm{kJ}}}}{{{\rm{mol}}}}} \right)} \right]\\\end{array} ΔHorxn=1mol×(986.1kJmol)[1mol×(635.1kJmol)+1mol×(285.8kJmol)]ΔHorxn=986.1kJ[635.1kJ285.8kJ]ΔHorxn=65.2kJ\begin{array}{c}\\\Delta {H^{\rm{o}}}_{{\rm{rxn}}} = 1{\rm{ }}\cancel{{{\rm{mol}}}} \times \left( { - 986.1{\rm{ }}\frac{{{\rm{kJ}}}}{{\cancel{{{\rm{mol}}}}}}} \right) - \left[ {1{\rm{ }}\cancel{{{\rm{mol}}}} \times \left( { - 635.1{\rm{ }}\frac{{{\rm{kJ}}}}{{\cancel{{{\rm{mol}}}}}}} \right) + 1{\rm{ }}\cancel{{{\rm{mol}}}} \times \left( { - {\rm{285}}{\rm{.8 }}\frac{{{\rm{kJ}}}}{{\cancel{{{\rm{mol}}}}}}} \right)} \right]\\\\\Delta {H^{\rm{o}}}_{{\rm{rxn}}} = - 986.1{\rm{ kJ}} - \left[ { - 635.1{\rm{ kJ}} - {\rm{285}}{\rm{.8 kJ}}} \right]\\\\\Delta {H^{\rm{o}}}_{{\rm{rxn}}} = - {\rm{65}}{\rm{.2 kJ}}\\\end{array}

And

ΔSSurroundings=ΔHorxnTΔSSurroundings=(65.2kJ)(25+273.15)KΔSSurroundings=(65.2kJ)298.15KΔSSurroundings=0.21868kJK\begin{array}{l}\\\Delta {S_{{\rm{Surroundings}}}} = - \frac{{\Delta {H^{\rm{o}}}_{{\rm{rxn}}}}}{T}\\\\\Delta {S_{{\rm{Surroundings}}}} = - \frac{{\left( { - {\rm{65}}{\rm{.2 kJ}}} \right)}}{{\left( {25 + 273.15} \right){\rm{ K}}}}\\\\\Delta {S_{{\rm{Surroundings}}}} = - \frac{{\left( { - {\rm{65}}{\rm{.2 kJ}}} \right)}}{{298.15{\rm{ K}}}}\\\\\Delta {S_{{\rm{Surroundings}}}} = 0.21868\frac{{{\rm{kJ}}}}{{\rm{K}}}\\\end{array}

The unit conversion is as follows:

1kJ=1000J1{\rm{ kJ}} = 1000{\rm{ J}}

ΔSSurroundings=0.21868kJK×1000J1kJΔSSurroundings=218.68JK\begin{array}{l}\\\Delta {S_{{\rm{Surroundings}}}} = 0.21868\frac{{\cancel{{{\rm{kJ}}}}}}{{\rm{K}}} \times \frac{{1000{\rm{ J}}}}{{1{\rm{ }}\cancel{{{\rm{kJ}}}}}}\\\\\Delta {S_{{\rm{Surroundings}}}} = 218.68\frac{{\rm{J}}}{{\rm{K}}}\\\end{array}

Calculate the value of ΔSUniverse\Delta {S_{{\rm{Universe}}}} as shown below.

ΔSUniverse=ΔSSystem+ΔSSurroundings=26.4JK+218.68JK=192.28JK\begin{array}{c}\\\Delta {S_{{\rm{Universe}}}} = {\rm{ }}\Delta {S_{{\rm{System}}}} + \Delta {S_{{\rm{Surroundings}}}}\\\\ = - 26.4{\rm{ }}\frac{{\rm{J}}}{{\rm{K}}}{\rm{ + 218}}{\rm{.68 }}\frac{{\rm{J}}}{{\rm{K}}}\\\\{\rm{ = 192}}{\rm{.28 }}\frac{{\rm{J}}}{{\rm{K}}}\\\end{array}

Thus, the value of ΔSUniverse\Delta {S_{{\rm{Universe}}}} is 192.28JK{\rm{192}}{\rm{.28 }}\frac{{\rm{J}}}{{\rm{K}}} .

Ans:

The value of ΔSUniverse\Delta {S_{{\rm{Universe}}}} for the reaction is 192.28JK{\rm{192}}{\rm{.28 }}\frac{{\rm{J}}}{{\rm{K}}} .